Wound Infection - Amikacin

JN, a 46-year old woman, has a wound infection. The culture report stated that the infection was caused by Pseudomonas aeruginosa. JN’s temperature was 102o F at the time of the culture. Amikacin sulfate is to be administered intravenously in 100 mL D5W over 45 minutes every 8 hours. The dosage is 15 mg/kg/day in three divided doses. JN weighs 165 pounds.

1. What is the drug classification of amikacin?

2. How many milligrams of amikacin should JN receive every 8 hours?

3. What type of IV equipment should be used for this infusion?

4. What would the IV infusion rate be?

5. When would a wound culture be obtained to determine the appropriate antibacterial agent? Explain your answer.

6. What are the similarities of amikacin to other aminoglycosides?

7. Wound one type of aminoglycoside be preferred over another? Explain your answer.

The nurse assesses JN for hearing and urinary function before and during the amikacin therapy.

8. What should a hearing assessment include?

9. JN’s urine output over the past 8 hours was 125 mL. Explain the possible cause for the amount of urine output. What nursing actions should be taken?

10. What laboratory tests monitor renal function?

11. The healthcare provider requests peak and trough serum amikacin levels. When should the blood samples be drawn to determine the peak serum level? The trough serum level?
Answer #1

1.Amikacin belongs to the group of antibiotics .It is classified under Aminoglycoside Antibiotics

2.Given

15mg/kg/day

Weight 165lbs (165*0.454=74.9 or 75kg)

As 1kg=15mg

Then 75kg=??mg

Cross multiply and divide

=75*15/1

=75*15

=1125mg/day

To calculate one dose divide 1125/3= 375mg

Ans:JN should receive 375mg of A Mika in every eight hours

3.The type of IV equipment which can he used are infusion pump in order to see it accurately or using the IV macroset tubing to infuse over 45minutes

4.Flow rate=Total volume (my) /Tine in minutes *Drop factor

Here volume=100ml

Time:45minutes

Drop factor:15gtt/ml (standard)

=(100/45)*15

=2.22*15

=33.3 or 33

Ans:The IV infusion rate will be 33drops/minute